OpenStudy (anonymous):

find sin(arccsc(20)),where arccsc(theta) is the inverse of csc(theta).

OpenStudy (anonymous):

So let \[\theta = arc\csc(20)\]We have \[\csc(\theta)=\frac{20}{1}\]Or in other words:\[\sin(\theta)=\frac{1}{20}\]Hmmmm.

OpenStudy (anonymous):

What happens when we replace \(\theta\) back with \(arc\csc(20)\)...

OpenStudy (anonymous):

@krypton Making sense?

OpenStudy (anonymous):

hm.. dont really know the topic,is there any rules for it?

OpenStudy (anonymous):

This is your own question.

OpenStudy (anonymous):

ok